If tanA=1\2 and tanB=1\3 then cos2A=
A sinB
B sin2B
C sin3B 
D none of these
 
 

Grade:11

2 Answers

Vikas TU
7 years ago
tan A =1/2
then 
cosA =2/root(5)
Cos2A= 2cos^2(A) -1 => 2*4/5 – 1 => 3/5
 
Now from the options,
Given tanB = 1/3
sinB = 1/root(10)
cosB = 3/root(10)
sin2B = 2sinBcosB = > 6/10 = 3/5
Thus C option.
Lalit kumar
6 years ago
Cos2A in terms of tan =1-tan2A/1+tan2A  after putting value we get cos2A= 3/5 now sin2A in tan form, sin 2B= 2tanB/1+tan2B put value of tanB we get 3/5